# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ Abstract
We consider bilevel optimization problems with general nonconvex lower-level objectives and show that the classical hyperfunction-based formulation is unsettled, since the global minimizer of the lower-level problem is generally unattainable. To address this issue, we propose a correspondence-driven hyperfunction $\phi^{\text{cd}}$. In this formulation, the follower is modeled not as a rational agent always attaining a global minimizer, but as an algorithm-based bounded rational agent whose decisions are produced by a fixed algorithm with initialization and step size. Since $\phi^{\text{cd}}$ is generally discontinuous, we apply Gaussian smoothing to obtain a smooth approximation $\phi^{\text{cd}}_{\xi}$, then show that its value and gradient converge to those of $\phi^{\text{cd}}$. In the nonconvex setting, we identify that bifurcation phenomena, which arise when $g(x,\cdot)$ has a degenerate stationary point, pose a key challenge for hyperfunction-based methods. This is especially the case when $\phi^{\text{cd}}_{\xi}$ is solved using gradient methods. To overcome this challenge, we analyze the geometric structure of the bifurcation set under some weak assumptions. Building on these results, we design a biased projected SGD-based algorithm SCiNBiO to solve $\phi^{\text{cd}}_{\xi}$ with a cubic-regularized Newton lower-level solver. We also provide convergence guarantees and oracle complexity bounds for the upper level. Finally, we connect bifurcation theory from dynamical systems to the bilevel setting and define the notion of fold bifurcation points in this setting. Under the assumption that all degenerate stationary points are fold bifurcation points, we analyze the Hessian behavior of the lower-level problem $g(x,\cdot)$ at its stationary points when the upper-level parameter $x$ lies in a neighborhood of the bifurcation set. We further characterize the manifold structure of the bifurcation set and establish the oracle complexity of SCiNBiO for the lower-level problem.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to\mathbb{R}$ is called the upper-level objective and $g:\mathbb{R}^{m}\to\mathbb{R}$ is called the lower-level objective; $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the feasible set for the upper-level problems. In game theory, the bilevel problem can be thought of as a two-player (Stackelberg) game, the lower-level and upper-level problems are also called the follower and the leader, respectively; see e.g. Kohli (2012); Liu et al. (2018).
When the lower-level problem satisfies certain regularity conditions, such as convexity or the Polyak-Łojasiewicz (PL) condition, two main approaches are commonly used to solve bilevel optimization problems: hypergradient-based methods (Ghadimi and Wang, 2018; Yang et al., 2021; Hong et al., 2023b) and value-function-based methods (Huang, 2023, 2024; Shen et al., 2025). Hypergradient-based methods aim to optimize the hyperfunction defined by
@@
where $\gamma$ is a penalty parameter. Similarly, to guarantee differentiability and computability of the value function $v(x)$, convexity or the PL assumption on the lower-level objective $g(x,y)$ is often imposed.
@@ 1.1 Related work
Lower-Level Global Optimality Guarantees: A notable nonconvex assumption enabling lower-level global optimality is the Polyak-Łojasiewicz (PL) condition, under which every stationary point is a global minimizer and gradient-based methods can find a global optimum. The works Huang (2023, 2024) propose hypergradient-based methods and prove non-asymptotic convergence under a PL condition for the lower-level problem; they also assume that the set of optimal solutions to the lower-level problem is a singleton which is restrictive. Shen et al. (2025) develops a value-function-based penalty method and establishes non-asymptotic convergence under a PL condition for the lower-level problem in $y$. In Chen et al. (2024), the authors design an algorithm that either requires the penalized hyperfunction to satisfy a PL condition in $y$, or assumes a PL condition in $y$ for the lower-level problem together with a singleton solution set, and proves non-asymptotic convergence under these assumptions. The study Chen et al. (2025) shows that, when the lower-level problem satisfies the PL condition in $y$, the optimistic hyper-objective function is weakly concave and the pessimistic hyper-objective function is weakly convex. A penalty-based approach is also taken by Jiang et al. (2025), who establish non-asymptotic convergence under the assumption that the penalized problem is PL in $y$. In Masiha et al. (2025), the authors assume a local PL condition (the PL inequality holds in a neighborhood of each local minimizer with a uniform constant) together with certain regularity conditions. Under these assumptions, they show that global optimality can still be attained. They propose a smoothing approximation of the pessimistic hyperfunction, design an algorithm, and establish its non-asymptotic convergence. Because minimax problems are special cases of bilevel optimization, it is relevant that Lu and Wang (2025) adopts a local KL assumption, weaker than the PL condition yet still implying global optimality, and under this assumption develops a proximal gradient-based method with non-asymptotic convergence.
@@
Without Lower-Level Global Optimality Guarantees: Many recent studies relax lower-level assumptions from guaranteeing global optimality to weaker ones, under which only a stationary point can be obtained. All of the following works fall into this category. The work Liu et al. (2022a) designs a barrier-based algorithm and, under a local PL assumption (PL condition holds locally around each local minimizer with a uniform constant), proves its non-asymptotic convergence. In Liu et al. (2024), a Moreau envelope-based reformulation is proposed, together with a penalty-based algorithm; under the assumption that the value function is weakly convex, the method is shown to converge to a stationary point of the Moreau envelope-based reformulation and achieves a non-asymptotic solution guarantee. The results in Bolte et al. (2025) show that if the lower-level problem is Morse in $y$ for any $x$, then there exist finitely many disjoint lower-level local-minimizer curves parameterized by $x$. A hyperfunction can thus be defined for each curve, and an algorithm is designed that converges asymptotically to a stationary point of one such hyperfunction. Xiao et al. (2025) proposes a hybrid algorithm that, without requiring strong assumptions, achieves asymptotic convergence to a KKT stationary point of the original problem. As a special case of bilevel optimization, the two-stage problem is also examined in Lou et al. (2025), where a constrained nonconvex lower-level problem is addressed via a log-barrier reformulation and an algorithm with asymptotic convergence guarantees is developed.
@@ 1.2 Contributions
  • We point out that some existing, relatively general assumptions still impose restrictions on the class of nonconvex problems, and introduce the notion of a prevalent assumption: one that holds with probability one after applying a specific type of arbitrarily small perturbation (e.g., adding a linear term) within a given function class. We prove that the assumption “for almost every $x$, the function $g(x,\cdot)$ is Morse” is prevalent (refer to Theorem 2.1), and identify that the set of $x$ for which $g(x,\cdot)$ is not Morse, although of measure zero, represents a key difficulty for hyperfunction-based algorithms in the nonconvex lower-level setting (see Section 2.22.3). We refer to this set as the bifurcation set (Definition 2.3);
@@ 2 Main difficulties
In this section, we discuss the main challenges in bilevel optimization with a nonconvex lower-level problem, affecting both problem formulation and algorithm design.
@@ 2.2 Limitations of Common Regularity Assumptions
Assume, for the sake of contradiction, that $\widetilde{g}(x,y)$ has no degenerate stationary points for any $x\in[0,1]$. Note that $\widetilde{g}(x,y)$ is continuously differentiable. By the implicit function theorem, if $\nabla_{y}\widetilde{g}(\bar{x},\bar{y})=0$ and the Hessian $\nabla^{2}_{yy}\widetilde{g}(\bar{x},\bar{y})$ is non-singular, then there exists a neighborhood $I$ of $\bar{x}$ and a unique smooth function $y:I\to\mathbb{R}$ with $y(\bar{x})=\bar{y}$ such that $\nabla_{y}\widetilde{g}(x,y(x))=0$ for every $x\in I$. Since $[0,1]$ is compact, we can cover it with finitely many such intervals $I_{1},...,I_{k}$ on each of which the implicit function theorem applies. Therefore, for any $\bar{x}\in[0,1]$ and stationary point $\bar{y}$ of $\widetilde{g}(\bar{x},\cdot)$, we can construct a smooth curve $x\mapsto y(x)$ such that $y(\bar{x})=\bar{y}$ and $\nabla_{y}\widetilde{g}(x,y(x))=0$ for every $x\in[0,1]$. Applying the above argument to the three non-stationary points of $\widetilde{g}(0,y)$, we obtain three smooth curves $y_{1}(x)$, $y_{2}(x)$, and $y_{3}(x)$ defined on $[0,1]$, such that for each $x\in[0,1]$, the point $y_{i}(x)$ is a stationary point of $\widetilde{g}(x,\cdot)$. By the local uniqueness guaranteed by the implicit function theorem, these curves cannot intersect, i.e., $y_{i}(x)\neq y_{j}(x)$ for $i\neq j$ and all $x\in[0,1]$. In particular, at $x=1$, the function $\widetilde{g}(1,y)$ must have at least three distinct stationary points. However, $\widetilde{g}(1,y)$ has only one stationary point. This contradiction implies that there must exist some $x\in(0,1)$ at which $\widetilde{g}(x,y)$ admits a degenerate stationary point.
@@
One can also consider the assumption that there exists a global constant $\mu>0$ such that $g(x,y)$ satisfies the local $\mu$-PL condition in $y$ for every $x$. This is a strong assumption, and the set of smooth functions satisfying it is not dense in the space of smooth functions. Consider again Example 2.1. At $x=0$, the function $g(0,y)$ has two non-degenerate local minimizers, whereas at $x=1$, $g(1,y)$ has one. This implies that as $x$ increase from 0 to $1$, at least one of the local minimizers of $g(0,y)$ must disappear through a degeneracy. Let $\bar{x}\in(0,1)$ denote the smallest point at which degeneration occurs. Consider the smooth function $y:[0,\bar{x})\to\mathbb{R}$ satisfying $\nabla_{y}g(x,y(x))=0$ for all $x\in[0,\bar{x})$, with initial value $y(0)$ equal to a non-degenerate local minimizer of $g(0,y)$. As $x$ approaches $\bar{x}$ from the left, the point $y(x)$ converges to a degenerate stationary point $y(\bar{x})$ of $g(\bar{x},y)$. For every $x\in[0,\bar{x})$, the point $y(x)$ is a locally strongly convex minimizer of $g(x,\cdot)$. As $x$ approaches $\bar{x}$ from the left, the local strong convexity constant at $y(x)$ tends to zero since $g(\bar{x},y)$ is degenerate at $y(\bar{x})$. Note that local strong convexity implies the local PL condition, and the PL constant $\mu(x)$ in this case coincides with the local strong convexity constant. It follows that $\mu(x)$ converges to 0 as $x$ approaches $\bar{x}$ from the left. Therefore, no uniform constant $\mu>0$ can exist such that $g(x,y)$ satisfies the $\mu$-PL condition in $y$ for all $x\in[0,1]$.
@@ 2.3 Beyond Pointwise Regularity: Prevalent Assumption and its Challenge
Let $\mathcal{V}$ be a topological vector space (e.g., the space of $C^{r}$ smooth functions), and let $\mathcal{S}\subseteq\mathcal{V}$ be a Borel-measurable subset. We say that $\mathcal{S}$ is prevalent if there exists a finite-dimensional subspace $P\subseteq\mathcal{V}$, called a probe set, such that for every $v\in\mathcal{V}$, the set $\{p\in P:v+p\in\mathcal{S}\}$ has full Lebesgue measure in $P$.
@@
Then, with probability one (i.e., almost surely with respect to the random choice of $a$), the bifurcation point set $\widetilde{\mathcal{X}}$ defined in (2.3) has measure zero.
@@
(ii) Parametric-family viewpoint. Now regard $g$ as a family $\{g(x,\cdot)\}_{x\in\mathcal{X}}$ of smooth functions, parametrised by $x$. A perturbation cannot simultaneously regularize every member of this family (see Example 2.1). Theorem 2.1 states that, after an arbitrarily small linear perturbation $g(x,y)+a^{\top}y$, which can be viewed as a uniform perturbation applied to the entire function family $\{g(x,\cdot)\}_{x\in\mathcal{X}}$, the probability (over $a$) that $g(x,\cdot)$ is Morse for almost every $x$ is 1. Thus, the condition
@@ 3 Correspondence-Driven Hyperfunction and Its Smoothing
We introduce in Section 3.1 a new definition of the hyperfunction called correspondence-driven hyperfunction based on algorithmic trajectories, which replaces the classical rational agent assumption of the lower level with a algorithm-based bounded rational agent. Unlike the classical rational agent, who is assumed to always attain a global minimizer even for nonconvex problems, an algorithm-based bounded rational agent generates its decision directly through the execution of a prescribed algorithm.

### v3

@@ Abstract
We consider bilevel optimization problems with general nonconvex lower-level objectives and show that the classical hyperfunction-based formulation is unsettled, since the global minimizer of the lower-level problem is generally unattainable. To address this issue, we propose a correspondence-driven hyperfunction $\phi^{\text{cd}}$. In this formulation, the follower is modeled not as a rational agent always attaining a global minimizer, but as an algorithm-based bounded rational agent whose decisions are produced by a fixed algorithm with initialization and step size. Since $\phi^{\text{cd}}$ is generally discontinuous, we apply Gaussian smoothing to obtain a smooth approximation $\phi^{\text{cd}}_{\xi}$, then show that its value and gradient converge to those of $\phi^{\text{cd}}$. In the nonconvex setting, we identify that bifurcation phenomena, which arise when $g(x,\cdot)$ has a degenerate stationary point, pose a key challenge for hyperfunction-based methods. This is especially the case when $\phi^{\text{cd}}_{\xi}$ is solved using gradient methods. To overcome this challenge, we analyze the geometric structure of the bifurcation set under some weak assumptions. Building on these results, we design a biased projected SGD-based algorithm SCiNBiO to solve $\phi^{\text{cd}}_{\xi}$ with a cubic-regularized Newton lower-level solver. We also provide convergence guarantees and oracle complexity bounds for the upper-level. Finally, we connect bifurcation theory from dynamical systems to the bilevel setting and define the notion of fold bifurcation points in this setting. Under the assumption that all degenerate stationary points are fold bifurcation points, we analyze the Hessian behavior of the lower-level problem $g(x,\cdot)$ at its stationary points when the upper-level parameter $x$ lies in a neighborhood of the bifurcation set. We further characterize the manifold structure of the bifurcation set and establish the oracle complexity of SCiNBiO for the lower-level problem.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to\mathbb{R}$ is called the upper-level objective and $g:\mathbb{R}^{m}\to\mathbb{R}$ is called the lower-level objective; $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the feasible set for the upper-level problem. In game theory, the bilevel problem can be thought of as a two-player (Stackelberg) game, the lower-level and upper-level problems are also called the follower and the leader, respectively; see e.g. Kohli (2012); Liu et al. (2018).
When the lower-level problem satisfies certain regularity conditions, such as convexity or the Polyak-Łojasiewicz (PL) condition, two main approaches are commonly used to solve bilevel optimization problems: hypergradient-based methods (Ghadimi and Wang, 2018; Yang et al., 2021; Hong et al., 2023b) and value-function-based methods (Huang, 2023, 2024; Shen et al., 2025; Lu and Mei, 2026). Hypergradient-based methods aim to optimize the hyperfunction defined by
@@
where $\gamma$ is a penalty parameter. Similarly, to guarantee differentiability and computability of the value function $v(x)$, strong convexity or the PL assumption on the lower-level objective $g(x,y)$ is often imposed.
@@ 1.1 Related work
Lower-Level Global Optimality Guarantees: A notable nonconvex assumption enabling lower-level global optimality is the Polyak-Łojasiewicz (PL) condition, under which every stationary point is a global minimizer and gradient-based methods can find a global optimum. The works Huang (2023, 2024) propose hypergradient-based methods and prove non-asymptotic convergence under a PL condition for the lower-level problem; they also assume that the set of optimal solutions to the lower-level problem is a singleton. Shen et al. (2025) develops a value-function-based penalty method and establishes non-asymptotic convergence under a PL condition for the lower-level problem in $y$. In Chen et al. (2024), the authors design an algorithm that either requires the penalized hyperfunction to satisfy a PL condition in $y$, or assumes a PL condition in $y$ for the lower-level problem together with a singleton solution set, and proves non-asymptotic convergence under these assumptions. The study Chen et al. (2025) shows that, when the lower-level problem satisfies the PL condition in $y$, the optimistic hyper-objective function is weakly concave and the pessimistic hyper-objective function is weakly convex. A penalty-based approach is also taken by Jiang et al. (2025), who establish non-asymptotic convergence under the assumption that the penalized problem is PL in $y$. In Masiha et al. (2025), the authors assume a local PL condition (the PL inequality holds in a neighborhood of each local minimizer with a uniform constant) together with certain regularity conditions. Under these assumptions, they show that global optimality can still be attained. They propose a smoothing approximation of the pessimistic hyperfunction, design an algorithm, and establish its non-asymptotic convergence. Because minimax problems are special cases of bilevel optimization, it is relevant that Lu and Wang (2025) adopts a local KL assumption, weaker than the PL condition, and under this assumption develops a proximal gradient-based method with non-asymptotic convergence.
@@
Without Lower-Level Global Optimality Guarantees: Many recent studies relax lower-level assumptions from guaranteeing global optimality to weaker ones, under which only a stationary point can be obtained. All of the following works fall into this category. The work Liu et al. (2022a) designs a barrier-based algorithm and, under a local PL assumption (PL condition holds locally around each local minimizer with a uniform constant), proves its non-asymptotic convergence. In Liu et al. (2024), a Moreau envelope-based reformulation is proposed, together with a penalty-based algorithm; under the assumption that the value function is weakly convex, the method is shown to converge to a stationary point of the Moreau envelope-based reformulation and achieves a non-asymptotic solution guarantee. In Lu (2025), the author introduces a stochastic primal-dual method based on a two-sided smoothing of the Moreau envelope reformulation; under the assumption that the lower-level objective is weakly convex, the algorithm is proven to converge to an $\epsilon$-KKT point with a non-asymptotic convergence. The results in Bolte et al. (2025) show that if the lower-level problem is Morse in $y$ for any $x$, then there exist finitely many disjoint lower-level local-minimizer curves parameterized by $x$. A hyperfunction can thus be defined for each curve, and an algorithm is designed that converges asymptotically to a stationary point of one such hyperfunction. Xiao et al. (2025) proposes a hybrid algorithm that, without requiring strong assumptions, achieves asymptotic convergence to a KKT stationary point of the original problem. As a special case of bilevel optimization, the two-stage problem is also examined in Lou et al. (2025), where a constrained nonconvex lower-level problem is addressed via a log-barrier reformulation and an algorithm with asymptotic convergence guarantees is developed.
@@ 1.2 Contributions
  • We point out that some existing, relatively general assumptions still impose restrictions on the class of nonconvex problems, and introduce the notion of a prevalent assumption: one that holds with probability one after applying a specific type of arbitrarily small perturbation (e.g., adding a linear term) within a given function class. We prove that the assumption “for almost every $x$, the function $g(x,\cdot)$ is Morse” is prevalent (refer to Theorem 2.1), and identify that the set of $x$ for which $g(x,\cdot)$ is not Morse, although of measure zero, represents a key difficulty for hyperfunction-based algorithms in the nonconvex lower-level setting (see Section 2.2, 2.3). We refer to this set as the bifurcation set (Definition 2.3);
@@
  • Through experiments, we demonstrate the effectiveness of SCiNBiO. First, we apply it to nonconvex-nonconcave minimax problems, which are a specific case of bilevel optimization with a nonconvex lower-level. The results show that SCiNBiO converges from all random initializations, whereas GDA struggles with cyclic behaviors and fails to converge in several cases. This highlights the stability and reliability of SCiNBiO in tackling challenging nonconvex problems. Second, we use SCiNBiO for bilevel hyperparameter optimization, specifically tuning a nonnegative $\ell_{2}$-regularization parameter $\lambda$ for a neural network. In this setting, SCiNBiO consistently outperforms V-PBGD and BOME in terms of both convergence speed and stability, demonstrating its superior performance in solving bilevel optimization problems with nonconvex lower-level objectives.
@@ 2 Main difficulties
In this section, we discuss the main difficulties in bilevel optimization with a nonconvex lower-level problem, affecting both problem formulation and algorithm design.
@@ 2.2 Limitations of Common Regularity Assumptions
Assume, for the sake of contradiction, that $\widetilde{g}(x,y)$ has no degenerate stationary points for any $x\in[0,1]$. Note that $\widetilde{g}(x,y)$ is continuously differentiable. By the implicit function theorem, if $\nabla_{y}\widetilde{g}(\bar{x},\bar{y})=0$ and the Hessian $\nabla^{2}_{yy}\widetilde{g}(\bar{x},\bar{y})$ is non-singular, then there exists a neighborhood $I$ of $\bar{x}$ and a unique smooth function $y:I\to\mathbb{R}$ with $y(\bar{x})=\bar{y}$ such that $\nabla_{y}\widetilde{g}(x,y(x))=0$ for every $x\in I$. Since $[0,1]$ is compact, we can cover it with finitely many such intervals $I_{1},...,I_{k}$ on each of which the implicit function theorem applies. Therefore, for any $\bar{x}\in[0,1]$ and stationary point $\bar{y}$ of $\widetilde{g}(\bar{x},\cdot)$, we can construct a smooth curve $x\mapsto y(x)$ such that $y(\bar{x})=\bar{y}$ and $\nabla_{y}\widetilde{g}(x,y(x))=0$ for every $x\in[0,1]$. Applying the above argument to the three non-degenerate stationary points of $\widetilde{g}(0,y)$, we obtain three smooth curves $y_{1}(x)$, $y_{2}(x)$, and $y_{3}(x)$ defined on $[0,1]$, such that for each $x\in[0,1]$, the point $y_{i}(x)$ is a stationary point of $\widetilde{g}(x,\cdot)$. By the local uniqueness guaranteed by the implicit function theorem, these curves cannot intersect, i.e., $y_{i}(x)\neq y_{j}(x)$ for $i\neq j$ and all $x\in[0,1]$. In particular, at $x=1$, the function $\widetilde{g}(1,y)$ must have at least three distinct stationary points. However, $\widetilde{g}(1,y)$ has only one stationary point. This contradiction implies that there must exist some $x\in(0,1)$ at which $\widetilde{g}(x,y)$ admits a degenerate stationary point.
@@
One can also consider the assumption that there exists a global constant $\mu>0$ such that $g(x,y)$ satisfies the local $\mu$-PL condition in $y$ for every $x$. This is a strong assumption, and the set of smooth functions satisfying it is not dense in the space of smooth functions. Consider again Example 2.1. At $x=0$, the function $g(0,y)$ has two non-degenerate local minimizers, whereas at $x=1$, $g(1,y)$ has one. This implies that as $x$ increases from 0 to $1$, at least one of the local minimizers of $g(0,y)$ must disappear through a degeneracy. Let $\bar{x}\in(0,1)$ denote the smallest point at which degeneration occurs. Consider the smooth function $y:[0,\bar{x})\to\mathbb{R}$ satisfying $\nabla_{y}g(x,y(x))=0$ for all $x\in[0,\bar{x})$, with initial value $y(0)$ equal to a non-degenerate local minimizer of $g(0,y)$. As $x$ approaches $\bar{x}$ from the left, the point $y(x)$ converges to a degenerate stationary point $y(\bar{x})$ of $g(\bar{x},y)$. For every $x\in[0,\bar{x})$, the point $y(x)$ is a locally strongly convex minimizer of $g(x,\cdot)$. As $x$ approaches $\bar{x}$ from the left, the local strong convexity constant at $y(x)$ tends to zero since $g(\bar{x},y)$ is degenerate at $y(\bar{x})$. Note that local strong convexity implies the local PL condition, and the PL constant $\mu(x)$ in this case coincides with the local strong convexity constant. It follows that $\mu(x)$ converges to 0 as $x$ approaches $\bar{x}$ from the left. Therefore, no uniform constant $\mu>0$ can exist such that $g(x,y)$ satisfies the $\mu$-PL condition in $y$ for all $x\in[0,1]$.
@@ 2.3 Beyond Pointwise Regularity: Prevalent Assumption and its Challenge
Let $\mathcal{V}$ be a topological vector space (e.g., the space of $C^{r}$ functions), and let $\mathcal{S}\subseteq\mathcal{V}$ be a Borel-measurable subset. We say that $\mathcal{S}$ is prevalent if there exists a finite-dimensional subspace $P\subseteq\mathcal{V}$, called a probe set, such that for every $v\in\mathcal{V}$, the set $\{p\in P:v+p\in\mathcal{S}\}$ has full Lebesgue measure in $P$.
@@
Then, with probability one (i.e., almost surely with respect to the random choice of $a$), $\widetilde{g}_{a}(x,y)$ is Morse in $y$ for almost every $x$.
@@
(ii) Parametric-family viewpoint. Now regard $g$ as a family $\{g(x,\cdot)\}_{x\in\mathcal{X}}$ of smooth functions, parameterized by $x$. A perturbation cannot simultaneously regularize every member of this family (see Example 2.1). Theorem 2.1 states that, after an arbitrarily small linear perturbation $g(x,y)+a^{\top}y$, which can be viewed as a uniform perturbation applied to the entire function family $\{g(x,\cdot)\}_{x\in\mathcal{X}}$, the probability (over $a$) that $g(x,\cdot)$ is Morse for almost every $x$ is 1. Thus, the condition
@@ 3 Correspondence-Driven Hyperfunction and Its Smoothing
We introduce in Section 3.1 a new definition of the hyperfunction called correspondence-driven hyperfunction based on algorithmic trajectories, which replaces the classical rational agent assumption of the lower-level with a algorithm-based bounded rational agent. Unlike the classical rational agent, who is assumed to always attain a global minimizer even for nonconvex problems, an algorithm-based bounded rational agent generates its decision directly through the execution of a prescribed algorithm.